OpenStudy (anonymous):

What is the length (magnitude) of the vector (2,-1)? a. √5 b. √2 c.-2 d.1

OpenStudy (anonymous):

( 2, -1) refers to the x y, we can form a triangle and use the pythaogoros theorom length = root ( x^2 + y^2) = root(2^2 + -1^2) = root ( 5 ) so the answer is A

OpenStudy (anonymous):

use \[a^2=b^2+c^2\] so \[a^2 = \pm\sqrt{2^2+1^1} = \pm\sqrt{5}\]

OpenStudy (anonymous):

so the answer is a